# AI-Driven Marine Robotics: Emerging Trends in Underwater Perception and Ecosystem Monitoring

### v1

@@ Abstract
Marine ecosystems face increasing pressure due to climate change, driving the need for scalable, AI-powered monitoring solutions. This paper examines the rapid emergence of underwater AI as a major research frontier and analyzes the factors that have transformed marine perception from a niche application into a catalyst for AI innovation. We identify three convergent drivers: environmental necessity for ecosystem-scale monitoring, democratization of underwater datasets through citizen science platforms, and researcher migration from saturated terrestrial computer vision domains. Our analysis reveals how unique underwater challenges—turbidity, cryptic species detection, expert annotation bottlenecks, and cross-ecosystem generalization—are driving fundamental advances in weakly supervised learning, open-set recognition, and robust perception under degraded conditions. We survey emerging trends in datasets, scene understanding and 3D reconstruction, highlighting the paradigm shift from passive observation toward AI-driven, targeted intervention capabilities. The paper demonstrates how underwater constraints are pushing the boundaries of foundation models, self-supervised learning, and perception, with methodological innovations that extend far beyond marine applications to benefit general computer vision, robotics, and environmental monitoring.
@@ 1 Introduction
This challenge has catalyzed the emergence of AI-driven underwater perception as a rapidly growing research frontier. The convergence of three key drivers is transforming what was once a niche application domain into a frontier for AI innovation (Fig. 1). First, the environmental imperative for scalable marine monitoring has created unprecedented demand for automated solutions that can operate at ecosystem scales. Second, the democratization of underwater data through citizen science initiatives and the release of large-scale datasets has provided the data foundation required for modern AI. Third, as traditional computer vision applications become increasingly saturated, researchers are turning to the underwater domain as a comparatively understudied environment that offers opportunities for novel methodological contributions while addressing pressing societal needs.
@@ 2 Why Now? Convergent Drivers
The emergence of AI for underwater perception and marine robotics as a major research frontier is not coincidental, but rather the result of three powerful forces converging at this moment in time. Understanding these drivers is crucial for appreciating both the urgency and the unprecedented opportunities that define this emerging field.
@@ 2.3 Research Opportunity
The underwater domain offers exactly this opportunity. Unlike terrestrial environments, underwater perception presents unique challenges that cannot be solved by simply applying existing methods: extreme lighting variations, backscatter and turbidity effects, color distortion with depth, and a diversity of marine life that exhibits camouflaging and cryptic behaviors. Foundation models developed for land-based applications typically do not directly translate for underwater applications, due to the sizable domain gap, largely caused by challenging underwater image characteristics, fine-grained classes and ecological considerations (Raine et al. 2024c), as explained further in Section 3. These challenges require genuine methodological innovation rather than incremental engineering improvements.
@@ 3 Emerging Trends
In this section, we provide a critical analysis of the recent and emerging trends in underwater perception. We do not provide an exhaustive survey of prior approaches, but instead highlight key works that we believe are driving innovation in the field and opening up new research directions. Where relevant, we also identify the specific challenges of the underwater domain driving the technical advances.
@@ 3.1 Datasets
One early initiative, the Virtual Reef Diver (2025), encouraged divers and snorkelers to upload reef images, which were then annotated by everyday citizens. These observations, combined with professional monitoring data, improved predictive maps of coral cover across the Great Barrier Reef (Vercelloni, Santos-Fernández, and Mengersen 2023). Similarly, Citizens of the Reef has shown that citizen scientists can reach expert-level accuracy in coral reef monitoring, with deep learning and crowd-sourced analysis producing 99% accurate benthic cover estimates across 8,086 images (Lawson et al. 2025). The Seatizen Atlas extends this model globally, contributing over 1.6 million geolocated images from the Southwest Indian Ocean collected by both citizens and researchers (Contini et al. 2025).
@@
Complementing these gamified initiatives, the Citizens of the Reef (2025) Great Reef Census integrates citizen science into education and corporate structures. The Schools Program allows students to learn about reef ecosystems while contributing data, and the Hour of Power initiative mobilizes CSR by encouraging employees from companies including Atlassian, Dell, Disney, Mars, Cotton On, and Salesforce to dedicate one hour to labeling reef imagery. Together, gamification and CSR expand participation, enhance ocean literacy, and strengthen stewardship of marine ecosystems.
@@ 3.2 Scene Understanding
The underwater environment presents a unique combination of perceptual and operational challenges that have historically limited automated scene understanding. However, these same constraints now serve as catalysts for AI innovation, driving the development of novel techniques that advance not only marine robotics but the broader fields of computer vision and machine learning. Rather than viewing underwater conditions as obstacles to overcome, the AI community is increasingly recognizing them as opportunities to push methodological boundaries in ways that benefit numerous application domains. As a result, underwater scene understanding has undergone a fundamental transformation, shifting from narrow task-specific models toward general-purpose foundation models that can handle the full complexity of marine ecosystems. This evolution mirrors broader trends in AI but has been accelerated by the unique demands of underwater environments.
@@
Marine ecosystems exhibit extraordinary biodiversity, with over 800 hard coral species alone (Dietzel et al. 2021), each requiring specialized taxonomic expertise to identify accurately (Fig. 2). Traditional supervised learning approaches to recognize marine species require massive datasets with expert annotations, a prohibitively expensive and slow process that creates a scalability bottleneck. This challenge has made the underwater domain a natural testbed for weakly supervised and few-shot learning approaches.
@@
Our recent point label aware superpixel method demonstrated that these sparse expert annotations can be effectively propagated by clustering pixels using learned features while respecting point label constraints, achieving substantial improvements while reducing computational requirements (Raine et al. 2022). This success has inspired even weaker supervision approaches that operate with image-level labels alone—requiring 25 times fewer annotations than patch-level methods (Raine et al. 2024b). For seagrass meadow monitoring, contrastive pre-training combined with feature similarity has enabled coarse segmentation from image-level supervision, outperforming traditional patch-supervised methods on multi-species classification tasks (Raine et al. 2024b).
The emergence of foundation models has further transformed this landscape, with DINOv2 features now capable of generating multi-species coral masks without any domain-specific pre-training (Raine et al. 2024c). When integrated with human-in-the-loop labeling strategies that leverage expert knowledge strategically, these approaches achieve remarkable efficiency gains (Raine et al. 2024c). This progression from dense supervision to sparse points to image-level labels, and finally to foundation model-assisted annotation, exemplifies how domain constraints can drive methodological innovation with broad applicability across fields requiring expert knowledge integration.
@@
There have been recent advances to tackle underwater visual degradation through transformer-based image enhancement (Chen et al. 2024; Qiao, Guan, and Zhi 2025). Sauder et al. (2023) introduced BackscatterNet and CausticsNet, self-supervised networks that estimate and correct backscatter and caustics from monocular sequences, improving real-time perception without ground-truth data. Similarly, physics-guided transformer architectures have been proposed to restore contrast and correct color shifts in challenging underwater conditions (Wang et al. 2023; Li, Chen, and Miao 2025).
@@
Marine organisms have evolved sophisticated camouflage and cryptic coloration strategies that represent the ultimate challenge for visual recognition systems (Fig. 3). Fish that perfectly mimic coral textures, octopi that dynamically change color and pattern, and invertebrates that are virtually indistinguishable from their substrates push the boundaries of what current computer vision can achieve (Hanlon 2007).
Recent investigations into performance disparities across marine species have uncovered that not all detection challenges are created equal, and that data quantity alone cannot explain why certain species consistently underperform in automated detection systems. Analysis of underwater object detection reveals that foreground-background discrimination represents the most critical bottleneck for cryptic species like scallops, regardless of training data availability. Even when class imbalance is artificially corrected, persistent precision gaps indicate that certain marine organisms present intrinsic feature-based challenges that cannot be resolved through data augmentation alone (Wille, Fischer, and Raine 2025). This finding has profound implications for computer vision more broadly, suggesting that some visual recognition problems may be fundamentally algorithmic rather than data-driven.
@@
Recent work has seen forays into applying vision–language models and large language models to the underwater domain, though this remains an ongoing area of research. For example, MarineGPT demonstrated the integration of language models into marine applications and highlighted the potential for multimodal understanding, where visual perception can be combined with domain-specific knowledge expressed in natural language (Zheng et al. 2023b). Similarly, natural language prompting with the Contrastive Language-Image Pretraining (CLIP) model (Radford et al. 2021) has been demonstrated for image pseudo-labeling, reducing annotation effort (Raine et al. 2024b). These studies highlight the early potential for using foundation models in specialized domains. MarineInst (Zheng et al. 2024a) builds on these ideas and is a foundation model for marine imagery which outputs both instance masks and captions for object instances.
@@ 3.3 Scene Reconstruction
Underwater 3D scene reconstruction has undergone a fundamental paradigm shift, moving from static photogrammetric workflows toward real-time neural rendering systems capable of dynamic ecosystem monitoring. This transformation represents a significant technical advance in marine robotics, with implications extending beyond underwater applications.
@@
The emergence of underwater-specific 3D Gaussian Splatting methods has substantially improved reconstruction capabilities, with systems like WaterSplatting (Li et al. 2025), SeaSplat (Yang, Leonard, and Girdhar 2025), and UW-GS (Wang et al. 2025) achieving real-time performance while incorporating physics-based underwater image formation models. These approaches explicitly model light attenuation, backscatter, and wavelength-dependent absorption; moving beyond generic computer vision techniques toward domain-aware algorithms that account for underwater optics.
@@
Learning-based Structure from Motion has enabled reconstruction of kilometer-length reef surveys processed in real-time (Sauder et al. 2024). Systems now achieve processing speeds that previously required hours in minutes, with semantic understanding integrated directly into the reconstruction pipeline. Recent advances in multi-view stereo combined with physical imaging models (Hu and Liu 2025) further accelerate processing while maintaining geometric accuracy. This represents a shift from post-hoc analysis to real-time ecosystem monitoring capabilities that can support active marine conservation efforts.
@@
The coupling of 3D reconstruction with semantic segmentation has progressed from technical demonstrations to operational marine science tools. Multi-modal approaches combining geometric reconstruction with automated species classification (Zhong et al. 2023; Remmers et al. 2025) enable direct computation of ecological metrics, transforming how marine biologists approach large-scale surveys and longitudinal studies. Gorry et al. (2025) have demonstrated that Visual Place Recognition can be integrated with feature matching and image segmentation to enable robust identification of revisited underwater areas and estimation of rigid transformations for ecosystem change analysis. These integrated systems demonstrate how AI-driven reconstruction can simultaneously provide geometric understanding and biological insights at scales previously impractical with traditional survey methods.
@@
Underwater reconstruction has evolved toward ecosystem-scale digital twins with demonstrated deployments spanning multi-million dollar research programs (Woods Hole Oceanographic Institution 2022) and into the tens to hundreds of hectares (King Abdullah University of Science and Technology 2025). These systems enable virtual experimentation with management interventions and provide the foundation for predictive ecosystem modeling at unprecedented scales. The development points toward reconstruction systems that capture dynamic, semantically-aware models supporting active marine conservation and real-time ecosystem management.
@@ 3.4 From Research to Impact
Robotics and AI are scaling coral propagation and rehabilitation efforts. A notable example is the Coral Growout Robotic Assessment System (CGRAS), which uses robotic imaging to automatically detect and count millions of baby corals in hatchery tanks (QUT 2024). This replaces thousands of hours of manual labor, providing scientists with real-time growth data and enabling large-scale coral deployment to the Great Barrier Reef.
@@
To ensure uptake, AI outputs are increasingly being mapped onto conventional monitoring protocols (e.g., percent cover from point-intercept). Platforms like ReefCloud and CoralNet operationalize random-point sampling and class predictions to yield comparable ecological indicators and dashboards (Australian Institute of Marine Science (2024) AIMS; Beijbom et al. 2012). For 3D products, standardized co-registration and uncertainty budgets allow statistically defensible temporal change estimates that align with regulatory expectations (Lechene et al. 2024).
@@ 4 Conclusions
AI for underwater perception illustrates a bidirectional relationship between environmental necessity and technological innovation. Conservation needs have accelerated research, while underwater constraints have driven advances in computer vision and machine learning that benefit many domains.
The convergent drivers, i.e. environmental imperative, data democratization, and research opportunity, have created a self-reinforcing ecosystem where new methods enable applications, generate data, and attract investment—transforming underwater AI into a recognized research frontier with real-world impact. Core challenges such as turbidity, cryptic species detection, and annotation bottlenecks have spurred innovations in weak supervision, open-set recognition, and robust perception in degraded environments. These advances extend beyond marine science to fields like medical imaging, satellite analysis, and autonomous systems.
@@
The dual impact identified here, AI enabling breakthroughs in marine science, while marine challenges advance AI methodology, shows how domain-specific engagement can drive both fundamental research and tangible impact.

### v2

@@ Abstract
Marine ecosystems face increasing pressure due to climate change, driving the need for scalable, AI-powered monitoring solutions to inform effective conservation and restoration efforts. This paper examines the rapid emergence of underwater AI as a major research frontier and analyzes the factors that have transformed marine perception from a niche application into a catalyst for AI innovation. We identify three convergent drivers: i) environmental necessity for ecosystem-scale monitoring, ii) democratization of underwater datasets through citizen science platforms, and iii) researcher migration from saturated terrestrial computer vision domains. Our analysis reveals how unique underwater challenges—turbidity, cryptic species detection, expert annotation bottlenecks, and cross-ecosystem generalization—are driving fundamental advances in weakly supervised learning, open-set recognition, and robust perception under degraded conditions. We survey emerging trends in datasets, scene understanding and 3D reconstruction, highlighting the paradigm shift from passive observation toward AI-driven, targeted intervention capabilities. The paper demonstrates how underwater constraints are pushing the boundaries of foundation models, self-supervised learning, and perception, with methodological innovations that extend far beyond marine applications to benefit general computer vision, robotics, and environmental monitoring.
@@ 1 Introduction
This bottleneck has catalyzed the emergence of AI-driven underwater perception as a rapidly growing research frontier. The convergence of three key drivers is transforming what was once a niche application domain into a frontier for AI innovation (Fig. 1). First, the environmental imperative for scalable marine monitoring has created unprecedented demand for automated solutions that can operate at ecosystem scales. Second, the democratization of underwater data through citizen science initiatives and the release of large-scale datasets has provided the data foundation required for modern AI. Third, as traditional computer vision applications become increasingly saturated, researchers are turning to the underwater domain as a comparatively understudied environment that offers opportunities for novel methodological contributions while addressing pressing societal needs.
@@ 2 Why Now? Convergent Drivers
The emergence of AI for underwater perception and marine robotics as a major research frontier is not coincidental, but rather the result of three powerful drivers converging at this moment in time. Understanding these drivers is crucial for appreciating both the urgency and the unprecedented opportunities that define this emerging field.
@@ 2.3 Research Opportunity
The underwater domain offers exactly this opportunity. Unlike terrestrial environments, underwater perception presents unique challenges that cannot be solved by simply applying existing methods: extreme lighting variations, backscatter and turbidity effects, color distortion with depth, and a diversity of marine life that exhibits camouflaging and cryptic behaviors. Foundation models developed for land-based applications typically do not directly translate to underwater applications, due to the sizable domain gap, largely caused by challenging underwater image characteristics, fine-grained classes, and ecological considerations (Raine et al. 2024c), as explained further in Section 3. These challenges require genuine methodological innovation rather than incremental engineering improvements.
@@ 3 Emerging Trends
This section provides a critical analysis of the recent and emerging trends in underwater perception. We do not provide an exhaustive survey of prior approaches, but instead highlight key works that we believe are driving innovation in the field and opening up new research directions. Where relevant, we also identify the specific challenges of the underwater domain driving the technical advances.
@@ 3.1 Datasets
One early initiative, the Virtual Reef Diver (2025), encouraged divers and snorkelers to upload reef images, which were then annotated by everyday citizens. These observations, combined with professional monitoring data, improved predictive maps of coral cover across the Great Barrier Reef (Vercelloni, Santos-Fernández, and Mengersen 2023). Similarly, Citizens of the Reef has shown that citizen scientists can reach expert-level accuracy in coral reef monitoring, with deep learning and crowd-sourced analysis producing 99% accurate benthic cover estimates across 8,086 images (Lawson et al. 2025). The Seatizen Atlas extends this model globally, contributing over 1.6 million geolocated images from the Southwest Indian Ocean collected by both citizens and researchers (Contini et al. 2025a).
@@
Complementing these gamified initiatives, the Citizens of the Reef (2025) Great Reef Census integrates citizen science into education and corporate structures. The Schools Program allows students to learn about reef ecosystems while contributing data, and the Hour of Power initiative mobilizes CSR by encouraging employees from companies including Atlassian, Dell, Disney, Mars, Cotton On, and Salesforce to contribute one hour of labor for labeling reef imagery. Together, gamification and CSR expand participation, enhance ocean literacy, and strengthen stewardship of marine ecosystems.
@@ 3.2 Scene Understanding
The underwater environment presents a unique combination of perceptual and operational challenges that have historically limited automated scene understanding. However, these same challenges now serve as catalysts for AI innovation, driving the development of novel techniques that advance not only marine robotics but the broader fields of computer vision and machine learning. Rather than viewing underwater conditions as obstacles to overcome, the AI community is increasingly recognizing them as opportunities to push methodological boundaries in ways that benefit numerous application domains. As a result, underwater scene understanding has undergone a fundamental transformation, shifting from narrow, task-specific models to general-purpose foundation models that can handle the full complexity of marine ecosystems. This evolution mirrors broader trends in AI but has been accelerated by the unique demands of underwater environments.
@@
Marine ecosystems exhibit extraordinary biodiversity, with over 800 hard coral species alone (Dietzel et al. 2021), each requiring specialized taxonomic expertise to identify accurately (Fig. 2). Traditional supervised learning approaches for marine species recognition rely on large, expertly-annotated datasets, a prohibitively expensive and slow process that creates a scalability bottleneck. This challenge has made the underwater domain a natural testbed for weakly supervised and few-shot learning approaches.
@@
Our recent point label aware superpixel method demonstrated that these sparse expert annotations can be effectively propagated by clustering pixels using learned features while respecting point label constraints, achieving substantial improvements while reducing computational requirements (Raine et al. 2022). This success has inspired even weaker supervision approaches that operate with image-level labels alone, requiring 25 times fewer annotations than patch-level methods (Raine et al. 2024b). For seagrass meadow monitoring, contrastive pre-training combined with feature similarity has enabled coarse segmentation from image-level supervision, outperforming traditional patch-supervised methods on multi-species classification tasks (Raine et al. 2024b).
The emergence of foundation models has further transformed this landscape, with DINOv2 features now capable of generating multi-species coral masks without any domain-specific pre-training (Raine et al. 2024c). When integrated with human-in-the-loop labeling strategies that strategically leverage expert knowledge, these approaches achieve remarkable gains in efficiency (Raine et al. 2024c). Similarly, Contini et al. (2025b) proposed a framework for weakly supervised semantic segmentation of coral reef aerial imagery, based on DINOv2 for generation of coarse spatial predictions and mask refinement with Segment Anything. This progression from dense supervision to sparse points to image-level labels, and finally to foundation model-assisted annotation, exemplifies how domain constraints can drive methodological innovation with broad applicability across fields requiring expert knowledge integration.
@@
There have been recent advances to tackle underwater visual degradation through transformer-based image enhancement (Chen et al. 2024; Qiao, Guan, and Zhi 2025). Sauder et al. (2023) introduced self-supervised networks that estimate and correct backscatter and caustics from monocular sequences, improving real-time perception without ground-truth data. Similarly, Wang et al. (2023) and Li, Chen, and Miao (2025) proposed physics-guided transformer architectures to restore contrast and correct color shifts in challenging underwater conditions.
@@
Marine organisms have evolved sophisticated camouflage and cryptic coloration strategies that represent the ultimate challenge for visual recognition systems (Fig. 3). Fish that perfectly mimic coral textures, octopuses that dynamically change color and pattern, and invertebrates that are virtually indistinguishable from their substrates push the boundaries of what current computer vision can achieve (Hanlon 2007).
Recent investigations into performance disparities across marine species have uncovered that not all detection challenges are created equal, and that data quantity alone cannot explain why certain species consistently underperform in automated detection systems. Analysis of underwater object detection reveals that foreground-background discrimination represents the most critical bottleneck for cryptic species like scallops, regardless of training data availability. Even when class imbalance is artificially corrected, persistent precision gaps indicate that certain marine organisms present intrinsic feature-based challenges that cannot be resolved through data augmentation alone (Wille, Fischer, and Raine 2026). This finding has profound implications for computer vision more broadly, suggesting that some visual recognition problems may be fundamentally algorithmic rather than data-driven.
@@
Recent work has seen forays into applying vision–language models and large language models to the underwater domain, though this remains an ongoing area of research. For example, MarineGPT demonstrated the integration of language models into marine applications and highlighted the potential for multimodal understanding, where visual perception can be combined with domain-specific knowledge expressed in natural language (Zheng et al. 2023b). Similarly, natural language prompting with the Contrastive Language-Image Pretraining (CLIP) model (Radford et al. 2021) has been demonstrated for image pseudo-labeling, reducing annotation effort (Raine et al. 2024b). MarineInst (Zheng et al. 2024a) builds on these ideas and is a foundation model for marine imagery which outputs both instance masks and captions for object instances, and UWBench (Zhang et al. 2025) proposes a comprehensive benchmark for underwater vision-language understanding. These studies highlight the early potential for using foundation models in specialized domains.
@@ 3.3 Scene Reconstruction
Underwater scene reconstruction has undergone a fundamental paradigm shift: from traditional photogrammetry towards high resolution 3D scene rendering systems, capable of producing photo-realistic digital twins for precise ecosystem monitoring. This transformation represents a significant technical advance in marine robotics, with implications extending beyond underwater applications.
@@
The emergence of underwater-specific 3D Gaussian Splatting methods has substantially improved reconstruction capabilities, with systems like WaterSplatting (Li et al. 2025), SeaSplat (Yang, Leonard, and Girdhar 2025), and UW-GS (Wang et al. 2025) achieving real-time rendering while incorporating physics-based underwater image formation models. These approaches explicitly model light attenuation, backscatter, and wavelength-dependent absorption; moving beyond generic computer vision techniques toward domain-aware algorithms that account for underwater optics.
@@
Learning-based Structure from Motion has enabled reconstruction of kilometer-length reef surveys in minutes, a process that previously required hours of manual labor (Sauder et al. 2024). These systems also integrate semantic segmentation directly into the reconstruction pipeline, unifying geometric mapping and scene understanding. Recent advances in multi-view stereo combined with physical imaging models (Hu and Liu 2025) further accelerate processing while maintaining geometric accuracy. This represents a shift from post-hoc analysis to rapid, in-situ ecosystem monitoring capabilities that can support active marine conservation efforts.
@@
The coupling of 3D reconstruction with semantic segmentation has progressed from technical demonstrations to operational marine science tools. Multi-modal approaches combining geometric reconstruction with automated species classification (Zhong et al. 2023; Remmers et al. 2025) enable direct computation of ecological metrics, transforming how marine biologists approach large-scale surveys and longitudinal studies. Gorry et al. (2025) have demonstrated that visual place recognition can be integrated with feature matching and image segmentation to enable robust identification of revisited underwater areas for ecosystem change analysis. These integrated systems demonstrate how AI-driven reconstruction can simultaneously provide geometric understanding and biological insights at scales previously impractical with traditional survey methods.
@@
Underwater reconstruction has evolved toward ecosystem-scale digital twins with demonstrated deployments spanning multi-million dollar research programs (Woods Hole Oceanographic Institution 2022) and into the tens to hundreds of hectares (King Abdullah University of Science and Technology 2025). These systems enable virtual experimentation with management interventions and provide the foundation for predictive ecosystem modeling at unprecedented scales. These underwater reconstruction systems comprise dynamic, semantically aware models which support active marine conservation and timely ecosystem management.
@@ 3.4 From Research to Impact
Robotics and AI are scaling coral propagation and rehabilitation efforts. A notable example is the Coral Growout Robotic Assessment System (CGRAS), which uses robotic imaging to automatically detect and count millions of baby corals in hatchery tanks (QUT 2024). This replaces thousands of hours of manual labor, providing scientists with real-time growth data and enabling large-scale coral deployment across the Great Barrier Reef.
@@
To ensure uptake, AI outputs are increasingly being mapped onto conventional monitoring protocols (e.g., percent cover from point-intercept). Platforms like ReefCloud and CoralNet operationalize random point sampling and class predictions to yield comparable ecological indicators and dashboards (Australian Institute of Marine Science (2024) AIMS; Beijbom et al. 2012). For 3D products, standardized co-registration and uncertainty budgets allow statistically defensible temporal change estimates that align with regulatory expectations (Lechene et al. 2024).
@@ 4 Conclusions
AI for underwater perception illustrates a bidirectional relationship between environmental necessity and technological innovation. Conservation needs have accelerated research, while underwater constraints have driven advances in computer vision and machine learning that benefit many domains. This dual impact demonstrates that domain-specific engagement can drive both fundamental research and tangible impact.
The convergent drivers, i.e. environmental imperative, data democratization, and research opportunity, have created a self-reinforcing ecosystem where new methods enable applications, generate data, and attract investment—transforming underwater AI into a recognized research frontier with real-world impact. Core challenges such as turbidity, cryptic species detection, and annotation bottlenecks have spurred innovations in weak supervision, open-set recognition, and robust perception in degraded environments. These advances extend beyond marine science to applications like medical imaging, analysis of satellite data, and autonomous systems.